The heights of male college students are normally distributed with



mean of 68 inches and standard deviation of 3 inches. If 80 samples consisting of 25 students each are drawn from the population, what would be the expected mean and standard deviation of the resulting sampling distribution of the means?

1
Expert's answer
2022-01-11T01:03:26-0500

The population is normally distributed with mean "68 \ in" and standard deviation "3\ in." There were a total of "80" samples of size "25." Average height was collected for each sample and these averages comprise the sampling distribution of means.

The mean of the sampling distribution is equal to the mean of the population


"\mu_{\bar{x}}=\mu=68\ in"

"N=80" is the population size, and "n=25" is the sample size.

Use the finite population correction (fpc) factor "\sqrt{\dfrac{N-n}{N-1}}"

If the population is of size "N ," if sampling is without replacement, and if the sample size is "n\leq N," then the standard error of the sampling distribution  is


"\sigma_{\bar{x}}=\dfrac{\sigma}{\sqrt{n}}\cdot\sqrt{\dfrac{N-n}{N-1}}"

"=\dfrac{3\ in}{\sqrt{25}}\cdot\sqrt{\dfrac{80-25}{80-1}}=0.5\ in"

The expected mean would be "68" inches, and standard deviation would be "0.5" inches.